A saturated solution of lead (II) iodide has an iodide ion concentration of 3.0 x 10-3. What is the molar solubility of PbI2? Determine the Ksp for PbI2.
Answer #1
PbI2 ===> Pb2+ + 2I- Ksp = [Pb2+][I-]^2 The molar solubility of PbCl2 = [Pb2+] Let [Pb2+] = x. Then [I-] = 2x, and (2x) = 3 * 10^-3 so x = 1.5 *10^-3 = molar solubility Ksp = x*x^2 = (1.5*10^-3)^3 = 3.375 * 10^-9
Answer #2
given [I-]=3x10^-3 as the molar ratio of Pb+2 an I- are in 2:1 ratio [Pb+2]=1.5x10^-3 Ksp=[Pb+2]*[I-]^2=1.5x9x10^-9=13.5x10^-9
Answer #3
PBI2=PB2+ + 2I- KSP=[Pb2+][I-]^2 =27*10^-9